Solve
x(x+3)- 2=2+x(x+1)
not in book

Answer by ivy12003(22)

x(x + 3) - 2 = 2 + x(x + 1)
x^2 + 3x - 2 = 2 + x^2 + x
3x - 2 = 2 + x
2x = 4
x = 2
To check, sub x = 2 into equation.
x(x + 3) - 2 = 2 + x(x + 1)
2(2 + 3) - 2 = 2 + 2(2 + 1)
2(5) - 2 = 2 + 2(3)
10 - 2 = 2 + 6
8 = 8